OpenStudy (lilcathyx):

If the sin 30° is 1/2, then the cos ____° = _____. A. 60° ; 1 over 2 B. 30° ; square root 2 over 2 C. 60° ; square root 3 over 2 D. 30° ; 1

OpenStudy (lilcathyx):

wouldn't it be "A"?

OpenStudy (anonymous):

Half of what?

OpenStudy (johnweldon1993):

Yes it would be A \[\large sin(x) = cos(90 - x)\] so \[\large sin(30) = cos(90 - 30) = cos(60)\]

OpenStudy (lilcathyx):

Thank you ! :O :D ^.^

OpenStudy (johnweldon1993):

haha anytime hun :)